# Continuous operators from spaces of Lipschitz functions

### v1

@@ 3. Density of spaces of Lipschitz functions
for all $x\in M$. Also, for each $n>0$ such that $G^{U,V}_{n}=\emptyset$, simply set $f^{U,V}_{n}(x)=0$ for all $x\in M$. Note that in either case $f^{U,V}_{n}$ is Lipschitz.
Since $\overline{G^{U,V}}\cap\overline{H^{U,V}}=\emptyset$, for every $x\in G^{U,V}$ there is $n>0$ such that $x\in G^{U,V}_{n}$. Also, for every $y\in H^{U,V}$ and $n\in\mathbb{N}$, it holds $\rho(y,G_{n}^{U,V})>0$. It follows that $G^{U,V}=\bigcup_{n=1}^{\infty}G^{U,V}_{n}$. Consequently, for every $x\in G^{U,V}$ and $y\in H^{U,V}$ there is $n>0$ such that $f^{U,V}_{n}(x)=0$ and $f^{U,V}_{n}(y)=1$, that is, $f^{U,V}_{n}$ separates $x$ and $y$.
@@
By the argument presented in the previous paragraph, $\mathcal{D}$ separates points, as for each $x\neq y\in M$ we can find $U,V\in\mathcal{B}$ such that $\varphi(x)\in U$, $\varphi(y)\in V$, and $\overline{U}\cap\overline{V}=\emptyset$. Hence, by the Stone–Weierstrass theorem for the compact-open topology (see [65, Problem 44B]), the algebra generated by $\mathcal{D}$ and the constant function $\textbf{1}\colon M\rightarrow\{1\}$ is dense in $C_{k}(M)$. Hence, the set $\mathcal{E}_{M}$ of all finite polynomial combinations, with rational coefficients, of members of $\mathcal{D}\cup\{\textbf{1}\}$ is dense in $C_{k}(M)$. Since $\mathcal{E}_{M}\subseteq\operatorname{LIP}(M)$, $\mathcal{E}_{M}$ is also dense in $(\operatorname{LIP}(M),\tau_{k})$. As $|\mathcal{E}_{M}|\leq|\mathcal{B}|=ww(M)$, we get that $d(\operatorname{LIP}(M),\tau_{k})\leq ww(M)$.
@@
We work with the topology $\tau$ on $C(M)$. The algebraic direct sum $C(M)=\mathbb{R}\oplus C_{e}(M)$ is topological (by [57, Corollary on p. 96]), therefore the canonical projection $P\colon C(M)\rightarrow C_{e}(M)$ is continuous. Consequently, $P$ transfers dense subsets of $C(M)$ onto dense subsets of $C_{e}(M)$. Since $P[\operatorname{LIP}(M)]=\operatorname{Lip}_{0}(M)$ and $\operatorname{LIP}(M)$ is dense in $C(M)$ (by Theorem 3.8.(1)), $\operatorname{Lip}_{0}(M)$ is dense in $C_{e}(M)$. Thus, (1) holds.
@@ 4. Continuous mappings from and onto spaces related to Lipschitz functions
Recall that, by a theorem of Grothendieck [31], for a compact space $K$ every uniformly bounded set in the Banach space $C(K)$ which is compact in $C_{p}(K)$ is weakly compact in $C(K)$. For any infinite metric space $M$ and its spaces $\operatorname{Lip}_{0}(M)$ and $\operatorname{Lip}_{0}(M)_{p}$ this theorem fails, as the unit ball $B_{\operatorname{Lip}_{0}(M)}$ of $\operatorname{Lip}_{0}(M)$ is compact in the pointwise topology (see [64, Theorem 2.37]) but not weakly compact (as $\operatorname{Lip}_{0}(M)$ is not reflexive). We also have the following negative result.
@@
The space $\operatorname{Lip}_{0}(M)$ is not weakly Lindelöf, since it contains a closed copy of the space $\ell_{\infty}$, which is well-known to be not weakly Lindelöf (see [18, Example 1.(i)]). Further, if $B_{\operatorname{Lip}_{0}(M)}$ was weakly Lindelöf, then $\operatorname{Lip}_{0}(M)_{w}$ would be covered by a countable collection of Lindelöf spaces, hence it would be itself a Lindelöf space. ∎
@@
By a similar argument as in (1), if there is a continuous surjection $\operatorname{Lip}_{0}(M)_{p}\rightarrow C_{p}(X)$, then $C_{p}(X)$ can be covered by a sequence of compact sets, which, by [5, Theorem I.2.1], means that $X$ is finite, which is a contradiction. Thus, (2) holds, too. ∎
@@
We showed in Corollary 3.4 that $d(\operatorname{Lip}_{0}(N)_{w})=d(\operatorname{Lip}_{0}(N))=2^{d(N)}$. Consequently we have, $d(\operatorname{Lip}_{0}(N),\tau^{\prime})\geq 2^{d(N)}$. On the other hand, we also have $d(C(X),\tau)=ww(X)\leq w(X)$ (by [52, Theorem 1], see the previous section), so $d(\operatorname{Lip}_{0}(N),\tau^{\prime})>d(C(X),\tau)$, hence there cannot exist any continuous mapping from $(C(X),\tau)$ onto $(\operatorname{Lip}_{0}(N),\tau^{\prime})$. ∎
Recall that $\operatorname{Lip}_{0}([0,1])$ is isomorphic to $\ell_{\infty}$, so in particular to the Banach space $(C(\beta\mathbb{N}),\|\cdot\|_{\infty})$. Since $w(\beta\mathbb{N})=2^{\aleph_{0}}=2^{d([0,1])}$, it follows that in Theorem 4.5 the assumption that $w(X)<2^{d(N)}$ cannot be relaxed to $w(X)\leq 2^{d(N)}$. On the other hand, Cúth, Doucha, and Wojtaszczyk [19, proof of Theorem 4.1] showed that there is no continuous linear surjection $T\colon C(K)\rightarrow\operatorname{Lip}_{0}([0,1]^{2})$ for any compact space $K$. Since the space $[0,1]^{2}$ is an absolute Lipschitz retract (see Section 5 for the definition), the latter result implies that for any metric space $M$ containing $[0,1]^{2}$ there is no continuous linear surjection $T\colon C(K)\rightarrow\operatorname{Lip}_{0}(M)$ for any compact space $K$ (as $\operatorname{Lip}_{0}([0,1]^{2})$ is complemented in $\operatorname{Lip}_{0}(M)$, see Proposition 5.5).
@@ 4.1. Continuous mappings from Lipschitz-free spaces
In this section we will briefly study the existence of continuous mappings from Lipschitz-free spaces $\mathcal{F}(M)$ (in particular, with respect to the weak topology).
Theorem 4.6.
@@
Let $M=C(K)$ for an uncountable compact metric space $K$. Let $N$ be an infinite compact ultrametric space (e.g. $N$ is the Cantor space $2^{\mathbb{N}}$). Then the Lipschitz-free Banach spaces $\mathcal{F}(M)$ and $\mathcal{F}(N)$ are homeomorphic but the spaces $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are not homeomorphic.
@@
Since $M$ is a separable Banach space, [29, Theorems 2.12 and 3.1] imply that $M$ embeds linearly into $\mathcal{F}(M)$ (as a closed subspace), so $\mathcal{F}(M)_{w}$ contains a linear copy of $C(K)_{w}$. Assume that $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are homeomorphic. Recall that $\mathcal{F}(N)$ is isomorphic to the space $\ell_{1}$ (by [22, Theorem 3.3]). By [27, Theorem 1.2] the space $(\ell_{1})_{w}$ is an $\aleph_{0}$-space. Consequently, $\mathcal{F}(M)_{w}$ is an $\aleph_{0}$-space, too. Since the property of being an $\aleph_{0}$-space is inherited by subspaces, the space $C(K)_{w}$ is also an $\aleph_{0}$-space, which is a contradiction to [51, Proposition 10.8]. ∎
@@
Suppose that (1) holds. Assume first that $d(M)>\aleph_{0}$ and $d(N)>\aleph_{0}$. Then, $M$ is not Lindelöf and hence by the Pytkeev theorem the tightness of $C_{p}(M)$ is uncountable (see [62, Exercise 149]). On the other hand, by the Kaplansky theorem (see [24, Theorem 3.54]), $\mathcal{F}(M)_{w}$ has countable tightness.
@@
Finally, assume that $d(M)>\aleph_{0}=d(N)$. Since $M$ is not separable, it is not second countable and hence $C_{p}(M)$ is not Lindelöf (see [62, Exercise 215]). Since $d(\mathcal{F}(N))=d(N)=\aleph_{0}$, $\mathcal{F}(N)$ is separable and hence $\mathcal{F}(N)_{w}$ is Lindelöf.
@@
If (3) holds, that is, $M$ is scattered, then by [26, Theorem 1.3] the space $C_{p}(M)$ is Ascoli. Suppose that $C_{p}(M)$ and $\mathcal{F}(N)_{w}$ are homeomorphic. Since the Ascoli property is inherited by continuous open maps (see [28, Proposition 3.3]), the space $\mathcal{F}(N)_{w}$ is also an Ascoli space. On the other hand, if the weak topology of a normed space $E$ is Ascoli, then $E$ is finite-dimensional (see [28, Proposition 3.2]), which is impossible in the case of $\mathcal{F}(N)$ for infinite $N$.
Assume that $M$ is not $\sigma$-compact, that is, (4) holds. By (1) we can assume that $\mathcal{F}(N)$ is separable, so, in particular, that it is a Polish space. Consequently, $\mathcal{F}(N)$ is analytic and hence $\mathcal{F}(N)_{w}$ is also analytic (as the identity is norm–weakly continuous). Were $C_{p}(M)$ and $\mathcal{F}(N)_{w}$ homeomorphic, $C_{p}(M)$ would be an analytic space as well. But, by [17], it would mean that $M$ is $\sigma$-compact, a contradiction.
Finally, assume that (5) or (6) holds. By (1) and (2) we only need to check the case when $|M|>\aleph_{0}$ and $d(\mathcal{F}(N))=\aleph_{0}$. But then, by [61, Proposition 4.4], $\mathcal{F}(N)_{w}$ is an $\aleph_{0}$-space, whereas, since $M$ is uncountable, $C_{p}(M)$ is not an $\aleph_{0}$-space, by [51, Proposition 10.7]. It follows that $C_{p}(M)$ and $\mathcal{F}(N)_{w}$ cannot be homeomorphic. ∎
We deal with Lipschitz-free Banach spaces having the Schur property in the next subsection. Note that, by [3, Theorem C], for no non-trivial Banach space $N$ the space $\mathcal{F}(N)$ can have the Schur property as in this case $N$ contains bilipschitz copies of subsets of the real line $\mathbb{R}$ of non-zero Lebesgue measure (e.g. of $\mathbb{R}$ itself). For Banach spaces $N$ and their spaces $\mathcal{F}(N)$ satisfying condition (6) of Theorem 4.8, we have the following proposition, related to the $\ell_{p}$-saturation. Recall that a Banach space $E$ is $\ell_{p}$-saturated for some $1\leq p<\infty$ if every infinite-dimensional closed linear subspace of $E$ contains an isomorphic copy of the space $\ell_{p}$.
Proposition 4.9.
Let $N$ be an infinite-dimensional separable Banach space such that $\mathcal{F}(N)$ isomorphically embeds into some space $L_{1}(\Omega,\Sigma,\mu)$. Then, every infinite-dimensional closed linear subspace of $N$ contains an isomorphic copy of the space $\ell_{q}$ for some $1\leq q\leq 2$. Moreover, if $\Omega$ is a topological space and $\Sigma$ is the Borel $\sigma$-field of $\Omega$, then every non-reflexive closed linear subspace of $N$ contains a complemented copy of the space $\ell_{1}$.
Proof.
Since $N$ is separable, [29, Theorem 3.1] implies that $N$ has the so-called lifting property and hence, by [29, Theorem 2.12], $N$ can be embedded as a closed linear subspace of the Lipschitz-free space $\mathcal{F}(N)$. Consequently, $N$ isomorphically embeds into $L_{1}(\Omega,\Sigma,\mu)$. It follows, by the theorem of Aldous [2, Theorem 1.1], that every infinite-dimensional closed linear subspace of $N$ contains an isomorphic copy of the Banach space $\ell_{q}$ for some $1\leq q\leq 2$.
For the “moreover” part, assume that $\Omega$ is a topological space and $\Sigma$ is the $\sigma$-field of all Borel subsets of $\Omega$. Again, by the lifting property, there is an isomorphic embedding $T\colon N\to L_{1}(\Omega,\Sigma,\mu)$. Then, for every non-reflexive closed linear subspace $P$ of $N$, the restriction $T|P$ is non-weakly compact, therefore, by the result of Pełczyński [53, Theorem 1], $P$ contains a complemented copy of $\ell_{1}$. ∎
@@ 4.2. The Schur property and sequential homeomorphisms
Taking into account that by [4, Corollary 2.7] the Lipschitz-free space on a discrete metric space has the Schur property, Theorem 4.10 implies the following characterization of the Schur property for Lipschitz-free spaces.
@@
    for every discrete metric space $N$ with cardinality $d(M)$, the spaces $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are sequentially homeomorphic.
@@
Assume that $\mathcal{F}(M)$ has the Schur property and let $N$ be a discrete metric space of cardinality $d(M)$. By the Toruńczyk theorem, the spaces $\mathcal{F}(M)$ and $\mathcal{F}(N)$ are homeomorphic, hence sequentially homeomorphic. Since both $\mathcal{F}(M)$ and $\mathcal{F}(N)$ have the Schur property, by Theorem 4.10, the spaces $\mathcal{F}(M)$ and $\mathcal{F}(M)_{w}$ are sequentially homeomorphic and the spaces $\mathcal{F}(N)$ and $\mathcal{F}(N)_{w}$ are sequentially homeomorphic. Now, (2) follows by composing appropriate sequential homeomorphisms.
The implication (2$\Rightarrow$(1) can be obtained by a similar argument. ∎

### v3

@@ 3. Density of spaces of Lipschitz functions
for all $x\in M$. Also, for each $n>0$ such that $G^{U,V}_{n}=\emptyset$, simply set $f^{U,V}_{n}(x)=0$ for all $x\in M$. Note that in either case $f^{U,V}_{n}$ is Lipschitz (see e.g. [18, Proposition 2.1.1]).
Since $\overline{G^{U,V}}\cap\overline{H^{U,V}}=\emptyset$, for every $x\in G^{U,V}$ there is $n>0$ such that $x\in G^{U,V}_{n}$ and hence $G^{U,V}=\bigcup_{n=1}^{\infty}G^{U,V}_{n}$. Consequently, for every $x\in G^{U,V}$ and $y\in H^{U,V}$ there is $n>0$ such that $f^{U,V}_{n}(x)=0$ and $f^{U,V}_{n}(y)=1$, that is, $f^{U,V}_{n}$ separates $x$ and $y$.
@@
By the argument presented in the previous paragraph, $\mathcal{D}$ separates points, as for each $x\neq y\in M$ we can find $U,V\in\mathcal{B}$ such that $\varphi(x)\in U$, $\varphi(y)\in V$, and $\overline{U}\cap\overline{V}=\emptyset$. Hence, by the Stone–Weierstrass theorem for the compact-open topology (see [67, Problem 44B]), the algebra generated by $\mathcal{D}$ and the constant function $\textbf{1}\colon M\rightarrow\{1\}$ is dense in $C_{k}(M)$. Hence, the set $\mathcal{E}_{M}$ of all finite polynomial combinations, with rational coefficients, of members of $\mathcal{D}\cup\{\textbf{1}\}$ is dense in $C_{k}(M)$. Since $\mathcal{E}_{M}\subseteq\operatorname{LIP}(M)$, $\mathcal{E}_{M}$ is also dense in $(\operatorname{LIP}(M),\tau_{k})$. As $|\mathcal{E}_{M}|\leq|\mathcal{B}|=ww(M)$, we get that $d(\operatorname{LIP}(M),\tau_{k})\leq ww(M)$.
@@
We work with the topology $\tau$ on $C(M)$. The algebraic direct sum $C(M)=\mathbb{R}\oplus C_{e}(M)$ is topological (by [59, Corollary on p. 96]), therefore the canonical projection $P\colon C(M)\rightarrow C_{e}(M)$ is continuous. Consequently, $P$ transfers dense subsets of $C(M)$ onto dense subsets of $C_{e}(M)$. Since $P[\operatorname{LIP}(M)]=\operatorname{Lip}_{0}(M)$ and $\operatorname{LIP}(M)$ is dense in $C(M)$ (by Theorem 3.8.(1)), $\operatorname{Lip}_{0}(M)$ is dense in $C_{e}(M)$. Thus, (1) holds.
@@ 4. Continuous mappings from and onto spaces related to Lipschitz functions
Recall that, by a theorem of Grothendieck [33], for a compact space $K$ every uniformly bounded set in the Banach space $C(K)$ which is compact in $C_{p}(K)$ is weakly compact in $C(K)$. For any infinite metric space $M$ and its spaces $\operatorname{Lip}_{0}(M)$ and $\operatorname{Lip}_{0}(M)_{p}$ this theorem fails, as the unit ball $B_{\operatorname{Lip}_{0}(M)}$ of $\operatorname{Lip}_{0}(M)$ is compact in the pointwise topology (see [66, Theorem 2.37]) but not weakly compact (as $\operatorname{Lip}_{0}(M)$ is not reflexive). We also have the following negative result.
@@
The space $\operatorname{Lip}_{0}(M)$ is not weakly Lindelöf, since it contains a closed copy of the space $\ell_{\infty}$, which is well-known to be not weakly Lindelöf (see [19, Example 1.(i)]). Further, if $B_{\operatorname{Lip}_{0}(M)}$ was weakly Lindelöf, then $\operatorname{Lip}_{0}(M)_{w}$ would be covered by a countable collection of Lindelöf spaces, hence it would be itself a Lindelöf space. ∎
@@
By a similar argument as in (1), if there is a continuous surjection $\operatorname{Lip}_{0}(M)_{p}\rightarrow C_{p}(X)$, then $C_{p}(X)$ can be covered by a sequence of compact sets, which, by [4, Theorem I.2.1], means that $X$ is finite, which is a contradiction. Thus, (2) holds, too. ∎
@@
We showed in Corollary 3.4 that $d(\operatorname{Lip}_{0}(N)_{w})=d(\operatorname{Lip}_{0}(N))=2^{d(N)}$. Consequently we have, $d(\operatorname{Lip}_{0}(N),\tau^{\prime})\geq 2^{d(N)}$. On the other hand, we also have $d(C(X),\tau)=ww(X)\leq w(X)$ (by [55, Theorem 1], see the previous section), so $d(\operatorname{Lip}_{0}(N),\tau^{\prime})>d(C(X),\tau)$, hence there cannot exist any continuous mapping from $(C(X),\tau)$ onto $(\operatorname{Lip}_{0}(N),\tau^{\prime})$. ∎
Recall that $\operatorname{Lip}_{0}([0,1])$ is isomorphic to $\ell_{\infty}$, so in particular to the Banach space $(C(\beta\mathbb{N}),\|\cdot\|_{\infty})$. Since $w(\beta\mathbb{N})=2^{\aleph_{0}}=2^{d([0,1])}$, it follows that in Theorem 4.5 the assumption that $w(X)<2^{d(N)}$ cannot be relaxed to $w(X)\leq 2^{d(N)}$. On the other hand, Cúth, Doucha, and Wojtaszczyk [21, proof of Theorem 4.1] showed that there is no continuous linear surjection $T\colon C(K)\rightarrow\operatorname{Lip}_{0}([0,1]^{2})$ for any compact space $K$. Since the space $[0,1]^{2}$ is an absolute Lipschitz retract (see Section 5 for the definition), the latter result implies that for any metric space $M$ containing $[0,1]^{2}$ there is no continuous linear surjection $T\colon C(K)\rightarrow\operatorname{Lip}_{0}(M)$ for any compact space $K$ (as $\operatorname{Lip}_{0}([0,1]^{2})$ is complemented in $\operatorname{Lip}_{0}(M)$, see Proposition 5.5).
@@ 4.1. Continuous mappings from Lipschitz-free spaces
In this section we will briefly study the existence of continuous mappings from Lipschitz-free spaces $\mathcal{F}(M)$ (in particular, with respect to the weak topology). We start with the following proposition.
Proposition 4.6.
@@
Let $M=C(K)$ for an uncountable compact metric space $K$. Let $N$ be a separable infinite ultrametric space (e.g. $N$ is the Cantor space $2^{\mathbb{N}}$). Then the Lipschitz-free Banach spaces $\mathcal{F}(M)$ and $\mathcal{F}(N)$ are homeomorphic but the spaces $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are not homeomorphic.
@@
Since $M$ is a separable Banach space, [31, Theorems 2.12 and 3.1] imply that $M$ embeds linearly into $\mathcal{F}(M)$ (as a closed subspace), so $\mathcal{F}(M)_{w}$ contains a linear copy of $C(K)_{w}$. Assume that $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are homeomorphic. Recall that $\mathcal{F}(N)$ is isomorphic to the space $\ell_{1}$ (by [20, Theorem 2]). By [29, Theorem 1.2] the space $(\ell_{1})_{w}$ is an $\aleph_{0}$-space. Consequently, $\mathcal{F}(M)_{w}$ is an $\aleph_{0}$-space, too. Since the property of being an $\aleph_{0}$-space is inherited by subspaces, the space $C(K)_{w}$ is also an $\aleph_{0}$-space, which is a contradiction to [54, Proposition 10.8]. ∎
@@
Suppose that (1) holds. Assume first that $d(M)>\aleph_{0}$ and $d(N)>\aleph_{0}$. Then, $M$ is not Lindelöf and hence by the Pytkeev theorem the tightness of $C_{p}(M)$ is uncountable (see [64, Exercise 149]). On the other hand, by the Kaplansky theorem (see [25, Theorem 3.54]), $\mathcal{F}(M)_{w}$ has countable tightness.
@@
Finally, assume that $d(M)>\aleph_{0}=d(N)$. Since $M$ is not separable, it is not second countable and hence $C_{p}(M)$ is not Lindelöf (see [64, Exercise 215]). Since $d(\mathcal{F}(N))=d(N)=\aleph_{0}$, $\mathcal{F}(N)$ is separable and hence $\mathcal{F}(N)_{w}$ is Lindelöf.
@@
If (3) holds, that is, $M$ is scattered, then by [28, Theorem 1.3] the space $C_{p}(M)$ is Ascoli. Suppose that $C_{p}(M)$ and $\mathcal{F}(N)_{w}$ are homeomorphic. Since the Ascoli property is inherited by continuous open maps (see [30, Proposition 3.3]), the space $\mathcal{F}(N)_{w}$ is also an Ascoli space. On the other hand, if the weak topology of a normed space $E$ is Ascoli, then $E$ is finite-dimensional (see [30, Proposition 3.2]), which is impossible in the case of $\mathcal{F}(N)$ for infinite $N$.
Assume that $M$ is not $\sigma$-compact, that is, (4) holds. By (1) we can assume that $\mathcal{F}(N)$ is separable, so, in particular, that it is a Polish space. Consequently, $\mathcal{F}(N)$ is analytic and hence $\mathcal{F}(N)_{w}$ is also analytic (as the identity is norm–weakly continuous). Were $C_{p}(M)$ and $\mathcal{F}(N)_{w}$ homeomorphic, $C_{p}(M)$ would be an analytic space as well. But, by [12, Theorem 2.1.3], it would mean that $M$ is $\sigma$-compact, a contradiction.
Finally, assume that (5) or (6) holds. By (1) and (2) we only need to check the case when $|M|>\aleph_{0}$ and $d(\mathcal{F}(N))=\aleph_{0}$. But then, by [63, Proposition 4.4], $\mathcal{F}(N)_{w}$ is an $\aleph_{0}$-space, whereas, since $M$ is uncountable, $C_{p}(M)$ is not an $\aleph_{0}$-space, by [54, Proposition 10.7]. It follows that $C_{p}(M)$ and $\mathcal{F}(N)_{w}$ cannot be homeomorphic. ∎
We deal with Lipschitz-free Banach spaces having the Schur property in the next subsection. Note that, by [2, Theorem C], for no non-trivial Banach space $N$ the space $\mathcal{F}(N)$ can have the Schur property as in this case $N$ contains bilipschitz copies of subsets of the real line $\mathbb{R}$ of non-zero Lebesgue measure (e.g. of $\mathbb{R}$ itself).
@@ 4.2. The Schur property and sequential homeomorphisms
Taking into account that by [3, Corollary 2.7] the Lipschitz-free space on a complete discrete metric space has the Schur property, Theorem 4.9 implies the following characterization of the Schur property for Lipschitz-free spaces.
@@
    for every complete discrete metric space $N$ with cardinality $d(M)$, the spaces $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are sequentially homeomorphic,
  3. (3)
    there is a complete discrete metric space $N$ with cardinality $d(M)$ for which the spaces $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are sequentially homeomorphic.
@@
Assume (1), that is, that $\mathcal{F}(M)$ has the Schur property and let $N$ be a complete discrete metric space of cardinality $d(M)$. By the Toruńczyk theorem, the spaces $\mathcal{F}(M)$ and $\mathcal{F}(N)$ are homeomorphic, hence sequentially homeomorphic. Since both $\mathcal{F}(M)$ and $\mathcal{F}(N)$ have the Schur property, by Theorem 4.9, the spaces $\mathcal{F}(M)$ and $\mathcal{F}(M)_{w}$ are sequentially homeomorphic and the spaces $\mathcal{F}(N)$ and $\mathcal{F}(N)_{w}$ are sequentially homeomorphic. Now, (2) follows by composing appropriate sequential homeomorphisms.
Implication (2$\Rightarrow$(3) is clear.
In order to show that (3) implies (1), let $N$ be a complete discrete metric space for which $\mathcal{F}(M)_{w}$ and $\mathcal{F}(N)_{w}$ are sequentially homeomorphic. Since $\mathcal{F}(N)$ has the Schur property, by Theorem 4.9 also the spaces $\mathcal{F}(N)$ and $\mathcal{F}(N)_{w}$ are sequentially homeomorphic. Since $\mathcal{F}(M)$ and $\mathcal{F}(N)$ have the same density, they are homeomophic by Toruńczyk’s theorem. Composing the corresponding mappings yields a sequential homeomorphism between $\mathcal{F}(M)_{w}$ and $\mathcal{F}(M)$, which, again by Theorem 4.9, yields that $\mathcal{F}(M)$ has the Schur property. ∎